OpenStudy (anonymous):

dy/dt = y(1-y)-hy Suppose that y(0) = y naught > 0 Solve IVP .

OpenStudy (anonymous):

separation of variables works...

OpenStudy (anonymous):

dy divides through .. but how does that look as integration

OpenStudy (anonymous):

\[\int\limits_{ }^{ } \frac{ dy }{ y(1-y-h) } = t+C\]

OpenStudy (anonymous):

I simplified a bit.

OpenStudy (anonymous):

could you possibly just show me where you simplified? thats part i'm having trouble with i think

OpenStudy (anonymous):

so we're just integrating the function (y(1-y-h))^-1

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

which is just ... ln|y(1-y-h)| ?

OpenStudy (anonymous):

over h-1?

OpenStudy (anonymous):

how's it coming?

OpenStudy (anonymous):

hmm... integral du/u = ln(u) but here, what would you pick for u? if you pick y(1-y-h), as you did, then du = ?

OpenStudy (anonymous):

wait why are we working with du now?

OpenStudy (anonymous):

hey!

OpenStudy (anonymous):

hey. got it yet?

OpenStudy (anonymous):

nooo!!! :( please help

zepdrix (zepdrix):

 Then from here, apply a Bournoulli Substitution, have you learned that yet? :o